884.One stick is 3 ft long and another is 6 ft long. You break the longer stick into sections. (a) If the sections are 2 ft and 4 ft long, will the sticks form a triangle?
(b) If the sections are 1 ft and 5 ft long, will the sticks form a triangle?
(c) If you break the longer stick at an arbitrary point, what is the probability that they form a triangle?

Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

One stick is 3 ft long and another is 6 ft long. You break the longer stick into sections.
(a) If the sections are 2 ft and 4 ft long, will the sticks form a triangle?
The longest side, 4 ft, must be longer than the difference between the lengths of the other 2 sides, and shorter than the sum of the other 2 sides.
===============================
(b) If the sections are 1 ft and 5 ft long, will the sticks form a triangle?
The longest side, 4 ft, must be longer than the difference between the lengths of the other 2 sides, and shorter than the sum of the other 2 sides.
=====================
(c) If you break the longer stick at an arbitrary point, what is the probability that they form a triangle?

Answer by greenestamps(13216)



The lengths of the two shorter sticks together must be greater than the length of the longest stick.

(a) The three sticks are now 2, 3, and 4 feet long. 2+3=5 is greater than 4; yes, you get a triangle.

(b) The three sticks are now 1, 3, and 5 feet long. 1+3=4 is less than 5; no, you don't get a triangle.

(c) Either with logical analysis, or with formal algebra, you can determine that the 6-foot stick can be cut at anywhere from 1.5 feet to 4.5 feet from one end to have the three sticks form a triangle (because the difference between 1.5 feet and 4.5 feet is 3 feet -- the length of the other stick.)

The range of where the cut can be made is then 4.5-1.5=3 feet; and the full length of the second stick is 6 feet. So the probability that the second stick gets cut in a way that makes it possible to form a triangle is 3/6 = 1/2.